OpenStudy (anonymous):

Find the indicated limit, if it exists. limit of f of x as x approaches 9 where f of x equals x plus 9 when x is less than 9 and f of x equals 9 minus x when x is greater than or equal to 9 The answers are: 18 0 9 The limit does not exist. I think the limit is 18. Am I right?

OpenStudy (freckles):

\[\lim_{x \rightarrow 9-}(x+9) \\ \lim_{x \rightarrow 9^+}(9-x)\]?

OpenStudy (anonymous):

This is how it looks

OpenStudy (freckles):

oh ok

OpenStudy (freckles):

we you need to find left and right limit

OpenStudy (freckles):

that is plug in 9 for both functions on the sides of x=9

OpenStudy (freckles):

if they equal, the the limit exist

OpenStudy (anonymous):

correct

OpenStudy (freckles):

if they don't equal, then the limit doesn't exist

OpenStudy (anonymous):

9 + 9 = 18 27 - 9 = 18 So the limit is 18? @freckles

OpenStudy (freckles):

sounds fab

OpenStudy (anonymous):

Could you help with another?

OpenStudy (anonymous):

Thank you by the way!

OpenStudy (freckles):

sure what is it

OpenStudy (anonymous):

The answers are 1.3 -1 4 5 I think it's 1.3. Is that right?

OpenStudy (anonymous):

Find limit as x approaches two from the left of f of x.

OpenStudy (freckles):

so you are saying 1.3 because it looks like that hols is above y=1

OpenStudy (freckles):

and yes that looks about right then

OpenStudy (freckles):

gj

OpenStudy (anonymous):

Basically.

OpenStudy (freckles):

 yeah basically we are only curious about the left hand side of x=2 (and also we don't even care what happens at x=2)